OpenStudy (anonymous):

Give the coordinates of two points on the line x = 5. Give the coordinates of two points on the line y = -2. Write the equation of the horizontal line passing through the point (-1, -8). Write the equation of the vertical line passing through the point (3, 1). Is the line with equation x = 4 horizontal or vertical? What is the slope of this line? Is the line with equation y = -8 horizontal or vertical? What is the slope of this line?

OpenStudy (anonymous):

There are six problems. (: Please help.. please (:

OpenStudy (anonymous):

Help please..

OpenStudy (anonymous):

1. (5,-1)

OpenStudy (anonymous):

and also (5,-2)

OpenStudy (anonymous):

It could be both? ?? :()

OpenStudy (mertsj):

the line x = 5 tells you that x MUST BE 5. It doesn't say anything about y so you get to pick anything for y that you please. Just make sure x is 5 as in (5,0) (5,3) (5,-90) and so on.

OpenStudy (anonymous):

oh okay thanks

OpenStudy (anonymous):

well I named two points that is on the line x=5 and there are many. So just write any point on that line:)

OpenStudy (mertsj):

Now for y = -2 the problem tells you that y MUST BE -2. It doesn't say anything about x so you get to pick whatever you want for x. Just make sure that y is -2 as in (5,-2) (-30,-2) (57,-2) and so on.

OpenStudy (anonymous):

1. you can just put any number for the y coordinate as long as the x coordinate is 5 2. the same for this one, as long as the y coordinate is -2 you would be right 3. y=-8 4.x=3 5.this line is vertical 6. this line is horizontal i think the slope for these lines would be zero

OpenStudy (mertsj):

If there is a horizontal line that passes through (-1,-8) the y value would always be -8 so that the line doesn't go up or down. So if y must be -8 we should write y = -8 and that is the equation.

OpenStudy (mertsj):

For a vertical line through (3,1), the line cannot go to the right or left so the x value has to stay right on 3. So write x = 3 and that is the equation.

OpenStudy (anonymous):

oh okay that makes sense

OpenStudy (mertsj):

The line x = 4 is vertical since it doesn't go to the right or left.

OpenStudy (mertsj):

And its slope is undefined.

OpenStudy (mertsj):

The slope y = -8 is horizontal and its slope is 0

OpenStudy (anonymous):

omg... thank u so much

OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (-6, 4) and (2, 0). Answer y - 4 = -one half (x + 6) y - 4 = 2(x + 6) y + 6 = -one half (x - 4) y + 6 = 2(x - 4)